OpenStudy (anonymous):

Solve: -20 - 3a = - 5(a - 2)

OpenStudy (anonymous):

\[-20 - 3a = -5(a - 2)\]\[\LARGE Distribute!\]\[-20 - 3a = -5a + 10\]Can you solve from here?

OpenStudy (pokemon23):

FOLLOW PEMDAS LEFT TO RIGHT IN THIS CASE THE GUY ABOVE ME SAID DISTRIBUTE.

OpenStudy (anonymous):

Simplifying further... \[-20 - 3a = -5a + 10 \implies 2a = 30 \implies a = ?\]

OpenStudy (anonymous):

thank u guys

OpenStudy (anonymous):

Glad to help :)

OpenStudy (pokemon23):

At your service.

OpenStudy (anonymous):

15

OpenStudy (pokemon23):

A=15 correct